# VisionUnite: A Vision-Language Foundation Model for Ophthalmology Enhanced with Clinical Knowledge

### v1

@@ Results
The architecture of VisionUnite is illustrated in Figure 1 (c), featuring a Transformer-based vision encoder, a Vision adapter, and a Large language model fine-tuned on LLaMA model [11]. Our innovative approach incorporates three distinct training objectives—visual text contrastive learning loss (CLIP Loss), visual classification loss (CLS Loss), and text generation loss (LLM Loss)—to facilitate the convergence of large-scale visual language models and enhance text generation quality. For the pre-training of VisionUnite, we construct a comprehensive dataset of 1.19 million pretrained image-text pairs including natural image-text pairs [12] and biomedical image-text pairs [13, 14, 15]. The architecture of VisionUnite incorporates three training objectives: image-text contrastive learning, classification supervised learning, and text-generation supervised learning. These objectives help refine vision encoders, accurately categorize features, and guide the final text output. During pre-training, we use text-generation supervised learning as the primary objective to build robust connections between images and corresponding texts. Furthermore, we construct a Multimodal Fundus fine-tuning dataset (MMFundus) to train and evaluate the diagnostic performance of VisionUnite. The MMFundus dataset is currently the largest multimodal fundus dataset, including 296,379 sets of fundus image-text pairs and corresponding constructed 889,137 rounds of dialogue. The MMFundus dataset covers the eight main categories of fundus image as shown in Figure 1 (b) and its image part includes 53 datasets, such as ODIR [16], APTOS [17], DeepDRiD [18], and REFUGE [19] dataset, etc. We design six sign categories including "Vascular", "Macular", "FBC (Fundus Boundary Color)", "OCD (Optical Cup Disc)", "FHE (Fundus Hemorrhages Examination)", and "Other" for each image. The six signs encompass all critical regions and functionalities of the eye, thereby providing a comprehensive assessment of eye health and disease. Each category targets a specific aspect of eye anatomy or pathology: Vascular for blood flow, Macular for central vision, FBC for boundary conditions, OCD for optic nerve health (glaucoma), FHE for retinal bleeding and exudates, and Other for miscellaneous conditions. We generate the corresponding sign labels for each image using the original label information of fundus images based on the above standards. More details of signs and dataset construction are in the methods section.
To assess the efficacy of models, we extract 180 samples and corresponding 540 rounds of questions from the MMFundus dataset to constitute the test dataset. It encapsulates a wide range of fundus conditions, ranging from health to illness status, encompassing Myopia, Macular Edema, Age-related Macular Degeneration, Glaucoma, Tessellation, Branch Retinal Vein Occlusion, Cataract, Optical Disc Pallor, Optical Disk Cupping, Drusen, Diabetic Retinopathy, Central Serous Retinopathy, and Arteriosclerotic Retinopathy, thirteen distinct ophthalmic diseases or conditions in total. We benchmark the performance of models against clinical diagnostics, focusing on two pivotal criteria: diagnostic accuracy and diagnostic relevance. Diagnostic accuracy gauges the precision of the diagnostic interpretations, encompassing disease identification and categorization. However, diagnostic accuracy cannot fully reflect the performance of large vision-language models in many cases. Different models may provide different explanations when giving the same diagnostic prediction, and we need relevant criteria to quantify this difference. Therefore, we design the diagnostic relevance, which provides insights into the response quality and further differentiates their differences. Our experimental framework involves presenting a fundus image and three related questions to a junior ophthalmologist (possessing one year of clinical experience). Concurrently, responses are elicited from three large vision-language models, 1. GPT-4V [20] (gpt-4-1106-vision-preview), 2. Gemini Pro [21] (gemini-pro-vision) and 3. our proposed VisionUnite. These sets of responses are then presented to three senior ophthalmologists (boasting over ten years of clinical experience), tasked with appraising the diagnostic accuracy and relevance. It is worth noting that we mandate the integration of responses into coherent paragraphs without extraneous AI-generated content to ensure an impartial evaluation. The details of diagnostic accuracy and relevance are as follows:
Diagnostic accuracy: We follow the clinical evaluation standards. 1. In cases of extra answers, we consider them to be incorrect answers. 2. In case of just missing unnecessary diagnostic information, we still consider the responses to be correct. 3. In our evaluation, the answer must only include all diagnosable diseases to be considered correct. We use the Wilson method [22] to estimate 95% confidence interval of diagnostic accuracy and calculate the p-value with the two-sided T-test and the above Wilson estimation.
Diagnostic relevance: 1. Senior ophthalmologists rank the response sets based on their alignment with the accurate diagnosis, scoring the most consistent response as four and the least as one. 2. Diagnostic relevance is designed to refine performance evaluation by taking into account factors that mere diagnostic accuracy might miss. For example, a response could be accurate in diagnostic prediction yet fail to adhere strictly to diagnostic criteria, which will diminish its overall diagnostic relevance. The bootstrap [23] method is used to estimate the 95% confidence interval of diagnostic relevance and the two-sided T-test is applied to calculate its P-value.
Additionally, we design a multiple-choice evaluation experiment to evaluate the diagnostic performance across various models objectively. Specifically, we provide four distinct options from all disease labels in the MMFundus dataset that models should select for the most likely diagnostic outcome based on the provided fundus image. The other three options are randomly selected except the correct options. We conduct experiments on 2233 cases and calculate the multiple-choice accuracy of ten diseases for two models: LLaVA-Med [24] (LLaVA-Med with finetuning) and VisionUnit. The results are in the appendix and Figure 11.
Comparison of Diagnosis between Ophthalmologist and Large Vision-Language Models
In our study, we conduct a comprehensive assessment of the diagnostic capabilities exhibited by the junior ophthalmologist compared with the large vision-language models, which include Gemini Pro, GPT-4V, and our proposed VisionUnite. The evaluation is performed on the designated test dataset. It encompasses a detailed analysis of diagnostic proficiency at both the disease and sign levels. Additionally, we explore an intricate examination of the diagnostic relevance between the responses provided by the junior ophthalmologist and those generated by the large vision-language models. As shown in Figure 2, the overall diagnostic accuracy of VisionUnite is over 45% and 28% higher than Gemini Pro and GPT-4V respectively, and the corresponding p-value is less than 0.001. The diagnostic relevance of VisionUnite is also over 62% and 25% higher than Gemini Pro and GPT-4V respectively with the p-value less than 0.001. Compared to the junior ophthalmologist, the overall accuracy of VisionUnite is approximately 4.5% higher with the p-value being 0.0876, and the diagnostic relevance is 0.75% higher (from 2.915 to 2.937). In each round of result evaluation, we find that the performance of VisionUnite in the first round of diagnostic Q&A is higher than that of junior ophthalmologist, with an accuracy rate of approximately 7.2% and a diagnostic relevance of 2.6%. The result indicates that VisionUnite has stronger analytical and reasoning abilities for fundus images, thus achieving superior performance in the first round of diagnosis. In the second and third rounds of diagnostic Q&A, the performance of VisionUnite is comparable to that of a junior ophthalmologist and far superior to GPT-4V and Gemini Pro. It also indicates that VisionUnite has the same problem-solving and interpretation abilities as doctors.
Diagnosis of Specific Disease or Condition and Sign
Building upon the foundational analysis of average diagnostic performance, our investigation further explores the nuanced diagnostic capabilities of the model and junior ophthalmologist for both healthy ocular conditions and a spectrum of specific ophthalmological pathologies. This segment of our study was particularly comprehensive, encompassing a wide array of conditions ranging from commonplace to rare. Specifically, our focus extended to seven distinct conditions including Age-Related Macular Degeneration, Branch Retinal Vein Occlusion, Cataract, Diabetic Retinopathy, Drusen, Glaucoma, and Myopia. This exhaustive approach allowed for a holistic understanding of the diagnostic acumen of these prevalent and less common ophthalmological diseases or conditions. In addition, we also explore the diagnostic performance involving different signs. As shown in Figure 3, VisionUnite and ophthalmologists perform significantly better than Gemini Pro and GPT-4V in six common and rare ophthalmic diseases or conditions except for Cataracts. The performance of GPT-4V on cataracts is comparable with VisionUnite and ophthalmologists. Among the diseases or conditions, VisionUnite has the best diagnostic performance in four diseases or conditions: Age-Related Macular Degeneration (AMD), Diabetic Retinopathy (DR), Drusen, and Glaucoma. Ophthalmologists perform better in three types of diseases or conditions: Branch Retinal Vein Occlusion, Cataract, and Myopia. For Cataract and Drusen, due to the small number of test samples, there may be significant differences in performance between different models. For AMD and DR, due to the tendency of ophthalmologists to provide incorrect disease staging, their performance is not as good as VisionUnite. Regarding diagnosis involving different signs, VisionUnite and ophthalmologists are also better than GPT-4V and Gemini Pro. As shown in Figure c1-c5, VisionUnite and ophthalmologists perform similarly on OCD (Optical Cup Disc). For FHE (Fundus Hemorrhages Examination) and Macular, VisionUnite performs better. For FBC (Fundus Boundary Color) and Vascular, ophthalmologists perform better. In terms of diagnostic relevance except for macular, ophthalmologists are better than all other models, which also reflects the ability of ophthalmologists to explain the results.
Joint Diagnosis of Multiple Diseases or Conditions and Signs
When dealing with fundus images from patients suffering from multiple diseases, the overlapping or combined manifestations of these conditions can significantly complicate the diagnostic process [7]. Each disease may present with specific signs on fundus examination, such as hemorrhages, exudates, or abnormalities in the vascular architecture. However, when multiple pathologies coexist, these signs can lead to a complex visual presentation that challenges even the most experienced ophthalmologists. For instance, diabetic retinopathy and arteriosclerotic Retinopathy both exhibit retinal vessel changes, but their specific impacts on the retina might slightly differ. Diabetic retinopathy typically shows more microaneurysms and hemorrhages, whereas arteriosclerotic retinopathy involves changes in the retinal arterioles, such as narrowing and tortuosity as shown in Figure 15. Therefore, we explore the joint diagnostic performance of the model and junior ophthalmologist for multiple diseases and multiple signs, including 15 diseases or conditions and 5 signs. As shown in Figure 7, the overall performance of VisionUnite is superior to GPT-4V and Gemini Pro in the presence of multiple diseases in fundus images. Compared to ophthalmologists, the overall accuracy of VisionUnite is similar, but the diagnostic relevance is lower. In the first two rounds, VisionUnite performs better than junior ophthalmologists, achieving a diagnostic accuracy of 44.44% and 77.78% with diagnostic relevance of 3.333 and 3.0. Gemini Pro performs the worst and cannot accurately make diagnosis predictions in the first round, with a diagnostic relevance of only 1.222. The results of multi-sign diagnosis are shown in Figure 7 d9-d16. Ophthalmologists perform best with 75% diagnostic accuracy and a diagnostic relevance of 3.375, followed by VisionUnite with 8.33% diagnostic accuracy of 5 and a diagnostic relevance of 2.917. The performance of GPT-4V and Gemini Pro is worse, with diagnostic accuracy of 35.42% and 12.5% respectively, and diagnostic relevance of 2.271 and 1.438. The results also reflect that the ability of models to analyze images is not as good as that of humans, making it difficult to obtain a correct diagnosis when encountering fundus images with multiple signs.
Misdiagnosis Analysis in Healthy Conditions
We also investigate the misdiagnosis of various models and ophthalmologists in the face of healthy fundus images. We use misdiagnosis rate and diagnostic relevance as evaluation indicators in this kind of scenario, where the misdiagnosis rate equals $1-diagnostic\ accuracy$, and diagnostic relevance remains as defined earlier. Since healthy samples often outnumber abnormal samples in the real clinical environment, which means it is important to evaluate the performance of the model and ophthalmologists on healthy samples. As shown in Figure 8, the overall misdiagnosis rate of VisionUnite is the lowest with only 8.33%, and its overall diagnostic relevance is 2.859. The ophthalmologists’ performance is at the second place with a misdiagnosis rate of 11.54% and a diagnostic relevance of 2.519. GPT-4V and Gemini Pro perform worse, with misdiagnosis rates of 19.87% and 42.95%, respectively. It is worth noting that in the first round of diagnosis, VisionUnite performed much better than other models and is also better than ophthalmologists, with a misdiagnosis rate of only 3.85% less than 15.38% of ophthalmologists, and with p-value is 0.05 compared with ophthalmologists. The results indicate that VisionUnite can extract good representations of healthy fundus images, thereby achieving a lower misdiagnosis rate through superior reasoning ability. However, the performance of the model is not as good as that of ophthalmologists in the final round involving the interpretation of fundus images and providing further opinions, which also indicates that the explanatory ability of the model needs to be further strengthened.
Diagnostic Correction in Patient Interaction
We assume that an advanced vision-language model ought to possess the capability for diagnostic rectification. It implies the inherent ability of models to autonomously amend inaccuracies in their initial diagnosis upon receiving additional information during a subsequent patient interaction. In this context, we focus on assessing the diagnostic precision of these models. Our evaluation encompassed a rigorous examination of the testing performance exhibited by VisionUnite, GPT-4V, and Gemini Pro. Notably, VisionUnite demonstrated superior corrective accuracy, which significantly surpassed GPT-4V and Gemini Pro. The superior performance of VisionUnite not only highlights the potential of domain-specific large models to exhibit heightened sensitivity in problem recognition within the medical field but also underscores their capacity for agile adaptation. Specifically, the ability of VisionUnite to swiftly recalibrate its responses in alignment with nuanced shifts in problem context, which we term as ’problem sensitivity’, stands as a testament to its refined diagnostic acumen and adaptability. As shown in Figure 9, we calculate the proportion of correct answers in the second or third round for each model in the case of incorrect answers in the first round. Overall refers to answering correctly in the second or third round. Round 2 refers to answering correctly in the second round, while Round 3 refers to answering correctly in the third round. The results demonstrate that VisionUnite has an overall correction accuracy of 86.67%, which is 24% and 39% higher than GPT-4V and Gemini Pro respectively with the p-value less than 0.001. In the second round, VisionUnite can perform much better than GPT-4V and Gemini Pro, with a second-round correction accuracy of 66.67% for VisionUnite, while GPT-4V and Gemini Pro can only achieve 28.04% and 20.69% respectively with the p-value also less than 0.001. We also evaluate the diagnostic accuracy of each model for both correct and incorrect predictions as shown in Figure 10. The findings demonstrate that VisionUnite’s performance in understanding problems and following instructions is equivalent to that of a junior ophthalmologist, independent of whether its predictions are correct or incorrect. The details can be seen in the appendix.
Diagnostic Errors Analysis between Ophthalmologist and Large Vision-Language Models
To analyze the outputs of the models thoroughly, we investigate their diagnostic errors and contrast them with assessments from ophthalmologists. We categorize these errors into two types: missed and incorrect errors. Missed errors pertain to incomplete yet correct diagnoses, while incorrect errors include entirely wrong diagnoses and partially correct diagnoses with additional, irrelevant errors. Our goal is to evaluate the completeness and accuracy of responses, noting omissions or irrelevant inclusions. Additionally, we assess error severity, which is categorized as error-free, minor, or major errors based on the impact on clinical judgment and treatment specificity. Minor errors encompass overlooked partial signs not critical to overall judgment or unnecessary diagnoses, as well as generalized treatment recommendations that lack specificity. Major errors include significantly incorrect responses. We also evaluate the potential for physical or mental harm resulting from the answers, grading potential health risks based on severity and likelihood. No harm is considered error-free, mild or moderate harm is considered minor errors, and serious harm or deaths are considered major errors. As shown in Figure 4, we analyze the diagnostic errors in five different situations: overall, single disease, multiple diseases, single sign, and multiple signs. Specifically, VisionUnite and ophthalmologists perform better than Gemini Pro and GPT-4V in various situations. Overall, VisionUnite performs the best in the "incorrect error" dimension, achieving an error-free rate of 78.15%, which is 4.82% higher than that of ophthalmologists with a corresponding p-value of 0.035. Ophthalmologists perform better in the "missed error" dimension, achieving an error-free rate of 71.85%. At the same time, we find that incorrect errors are more associated with major errors than minor errors. Compared to missed errors, incorrect errors often lead to more serious consequences. However, the results also indicate that there is still a certain gap between VisionUnite and ophthalmologists in the diagnosis of multiple diseases, and it also indicates that the understanding and reasoning abilities of models for images need to be further improved to reach the level of ophthalmologists. The details of the classification criteria for diagnostic errors are in Table 2 in the appendix.
Diagnostic analysis of VisionUnite assisted junior ophthalmologist
To further explore the clinical decision-support capabilities of VisionUnite, we assess its impact on the diagnostic performance of primary ophthalmologists. We specifically evaluate the improvements in diagnostic accuracy and efficiency when VisionUnite assists ophthalmologists in the initial round which includes 180 questions. Our analysis shows that VisionUnite facilitates a 26.98% reduction in diagnostic time and a 29.44% increase in diagnostic accuracy overall illustrated in Figure 5. Notably, the use of VisionUnite leads to a remarkable 50% improvement in diagnostic accuracy and a 33% reduction in diagnostic time in cases of diabetic retinopathy. Furthermore, we examine the impact of VisionUnite on diagnoses involving five distinct physical signs detailed in Figure 5 i8-i12. The assistance is most significant for fundus hemorrhage exudates (FHE) and optic disc cupping (OCD), where VisionUnite helps improve diagnostic accuracy by 39.02% and 35.49%, respectively, while also reducing diagnostic time by 35.63% and 24.25%. In traditional fundus lesion screening, the reliance on the subjective judgment of ophthalmologists based on fundus photography can lead to variability in diagnosis owing to human factors such as fatigue, variability in training, and personal interpretative skills. This subjectivity can contribute to higher risks of missed diagnoses and misdiagnoses. In contrast, VisionUnite can standardize the interpretation of fundus images by consistently applying the same criteria to analyze and interpret data across all instances. In diseases with multifaceted presentations, such as diabetic retinopathy, VisionUnite can analyze multiple aspects of the disease, providing a holistic view that aids in a more comprehensive assessment. Additionally, The predictive diagnostics of VisionUnite further assist physicians in developing and evaluating treatment strategies.
Consistent Interpretation of Visual and Language Features
In our final analytical phase, we scrutinize the coherent integration of visual and textual elements within large vision-language models, an endeavor aimed at assessing the alignment between the textual descriptions and the corresponding visual data. Our investigation addresses the potential for ’illusory discrepancies’ within the output of these sophisticated models. Specifically, we list the outputs of VisionUnite against those from other large vision-language models, with a keen focus on discerning any textual outputs that may depict features not present within the associated imagery. For instance, in cases where the diagnosis pertained to age-related macular degeneration, and the imagery exclusively showcases the presence of drusen, we should evaluate whether the textual narrative unjustifiably extended the diagnosis to encompass additional manifestations such as hemorrhaging and exudation. Such a detailed examination is pivotal for understanding and mitigating the tendency of these models to over-interpret or misalign textual descriptions with their visual counterparts, thereby ensuring a more accurate and reliable diagnostic output. Meanwhile, we also present a series of evaluation criteria for the responses of the models and junior ophthalmologists in this section. More cases from Junior Ophthalmologist and Large Vision-Language Models including VisionUnite, GPT-4V, and Gemini Pro can be seen in the extended data Figure 12-15 and the corresponding expert evaluation is in Figure 16.
Discussion
The diagnostic performance of VisionUnite in ophthalmic diseases has undergone thorough validation across diverse clinical scenarios, encompassing open-ended multiple disease diagnoses, clinical interpretation, and interactive inquiry from the patients. The integration of VisionUnite is poised to significantly elevate the precision of initial screenings for ophthalmic diseases, consequently enhancing the diagnostic efficiency of healthcare professionals, which is a particularly pivotal advancement in underdeveloped regions or those plagued by scarce ophthalmic services. Experimental findings underscore that VisionUnite’s capabilities in image analysis, disease diagnosis, and clinical explanations have achieved parity with ophthalmic junior doctors and, in certain instances, even surpassed their performance in diagnosing specific diseases. As attested by ophthalmologists, VisionUnite outshines general large vision-language models like GPT-4V and Gemini Pro, not only in delivering more accurate disease diagnoses but also in furnishing clear diagnostic foundations, which is an aspect often lacking in typical large vision-language models. In specific clinical applications, VisionUnite proves adaptable for patient inquiries while concurrently assisting doctors in optimizing diagnoses and facilitating ophthalmic training. Besides, VisionUnite expedites doctors’ disease diagnosis, offers question-and-answer services for patients to address their queries, and delivers tailored responses. For healthcare professionals, VisionUnite serves as a valuable auxiliary diagnostic tool, leveraging extensive ophthalmic knowledge and adeptness in disease diagnosis. Different from CT or X-ray, fundus photography typically lacks accompanying textual descriptions. Traditionally, the patient’s understanding of fundus imaging results primarily derives from doctors’ verbal explanations, which can sometimes lead to inadvertent omissions. However, the introduction of VisionUnite offers a robust solution by generating comprehensive fundus photography reports. The integration of verified VisionUnite reports not only aids doctors in the diagnostic process but also mitigates the risk of potential oversights.
While VisionUnite demonstrates diagnostic capabilities comparable to those of a junior ophthalmologist, certain limitations warrant consideration. Notably, the clinical interpretation of fundus images by VisionUnite can at times be overly broad, hindering the delivery of more precise information and consequently constraining performance enhancements in subsequent second-rounded diagnoses. Furthermore, the current proficiency of VisionUnite is confined to processing fundus images due to constraints in available data modalities. Future endeavors aim to expand its capabilities to include additional ophthalmic modalities such as FA, OCT, OCTA, and others. We also believe that the augmentation of data and modalities during the training process will enhance VisionUnite’s ability to handle diverse modalities and bolster its diagnostic performance. Another limitation lies in the current granularity of sign-level labels, as we have only categorized fundus features into six sign labels. It has led to omissions in predictions for images exhibiting multiple sign labels. To address the limitation, we plan to involve more fine-grained sign-level labels to better discern between various fundus images and enhance diagnostic accuracy. As computational power advances and more high-quality annotated data becomes available, we aspire for VisionUnite to extend its diagnostic reach to encompass a broader spectrum of rare diseases. Simultaneously, we envision its role as a valuable tool for ophthalmologists, aiding in identifying and classifying lesions and diseases. By introducing a more extensive array of sign-level labels, we also believe VisionUnite has the potential to unveil previously elusive links between signs and diseases. In addition, it could contribute to the identification of novel diagnostic evidence, fostering the evolution of diagnostic standards and protocols. Lastly, we expect VisionUnite to unveil connections between ocular features and systemic diseases. Contemporary research increasingly proposes the utility of ocular examinations for early screening of systemic conditions, such as Alzheimer’s disease. Leveraging the heightened visual encoding capabilities of large vision-language models and incorporating data on systemic diseases, we envision VisionUnite bridging ophthalmic data with information from other body regions. It is poised to expedite the revelation of hitherto undiscovered connections between ocular features and diseases manifesting in other organs.
Methods

### v2

@@ III-A Overview
The architecture of VisionUnite is illustrated in Figure 2 (a), featuring four key components: a Transformer-based vision encoder, a Vision adapter, a vision projector and a Large language model fine-tuned on LLaMA model [11]. For the pre-training of VisionUnite, we construct a comprehensive dataset of 1.19 million pretrained image-text pairs including natural image-text pairs [14] and biomedical image-text pairs [15, 16, 17]. The architecture of VisionUnite incorporates three training objectives: image-text contrastive learning, classification supervised learning, and text-generation supervised learning. These objectives help refine vision encoders, accurately categorize features, and guide the final text output. During pre-training, we use text-generation supervised learning as the primary objective to build robust connections between images and corresponding texts. Furthermore, we construct a Multimodal Fundus fine-tuning dataset (MMFundus) to train and evaluate the diagnostic performance of VisionUnite. The MMFundus dataset is currently the largest multimodal fundus dataset, including 296,379 sets of fundus image-text pairs and corresponding constructed 889,137 rounds of dialogue. The MMFundus dataset covers the eight main categories of fundus image as shown in Figure 1 (b) and its image part includes 53 datasets, such as ODIR [18], APTOS [19], DeepDRiD [20], and REFUGE [21] dataset, etc. We design six sign categories including ”Vascular”, ”Macular”, ”FBC (Fundus Boundary Color)”, ”OCD (Optical Cup Disc)”, ”FHE (Fundus Hemorrhages Examination)”, and ”Other” for each image. The six signs encompass all critical regions and functionalities of the eye, thereby providing a comprehensive assessment of eye health and disease.
III-B Datasets Construction
III-B1 Datasets for pre-training VisionUnite
In the pre-training phase of VisionUnite, we utilize a dataset comprising 1.24 million image-text pairs, encompassing around 616,435 natural image-text pairs sourced from the COCO dataset [14] and around 623,816 biomedical image-text pairs sourced from the PMC-OA dataset [15, 16] (608,027 images) and Retina Image Bank datasets [17] (15,789 images). The expansive PMC-OA dataset encapsulates over 1.6 million image-text pairs drawn from 25 million scientific literature images and titles. Rigorous screening and preprocessing efforts are applied to distill this dataset down to 623,816 image-text pairs as outlined in the appendix. Throughout the pre-training process, we just employ text-generation supervised learning as the pre-training objective. This kind of strategy aims to establish robust representation connections between images and their corresponding texts. In the context of dialogue, the answer section is represented by captions from the text-image pairs. We have constructed 20 questions for the question section, which are categorized into two types: indicative long answers (10 sentences) and short answers (10 sentences). Depending on the length of the dialogue answer, a sentence from the relevant question type is selected as the interrogative component of the dialogue. The comprehensive list of the 20 constructed questions is provided in the appendix for reference.
III-B2 MMFundus Dataset for fine-tuning VisionUnite
To realize the open-ended multi-disease diagnosis and comprehensive clinical explanations grounded in multiple rounds, we curate the pioneering multimodal fundus dataset, MMFundus. MMFundus dataset comprises 296,379 pairs of fundus images and corresponding text, accompanied by 889,137 rounds of dialogue. While VisionUnite leverages substantial datasets for training, we develop a semi-automated annotation framework requiring minimal expert supervision. This framework employs a three-stage verification protocol: 1. Initial rule-based labeling leveraging structured report metadata. 2. Automated label propagation through classification by visual similarity. 3. Selective expert verification of boundary cases (5% of the dataset). This methodology achieved 99% annotation accuracy while reducing expert labeling needs by 95% through an iterative self-training approach. Drawing from 52 public datasets and one private dataset, the image data in MMFundus embodies a diverse spectrum of ophthalmic conditions. The textual and dialogue components are sourced through doctor annotations, inherent category labels, and the automated generation capabilities of large language models such as InternLM [22] and GPT-4 [23], as the details are provided in the appendix. For the testset, all image descriptions and dialogue data have undergone annotation or confirmation by medical professionals. In constructing the training set, we initially utilized doctor annotations and category labels to establish a prototype dataset. Leveraging the prototype dataset alongside a large language model, we automatically generate the image-related descriptions and dialogues batch by batch. We also design a hierarchical approach to generate descriptions at three levels—Normal/Abnormal, disease name, and clinical explanation based on the disease within the dataset. Our method for ensuring the quality and accuracy of automatically generated content in the MMFundus dataset employs a comprehensive, multi-layered validation framework: (1) the expert-annotated prototype dataset establishing foundational parameters; (2) the hierarchical three-level annotation architecture with sign-level constraints that intrinsically validates semantic consistency; (3) an iterative refinement protocol where identified errors trigger complete batch regeneration until reaching 99% accuracy. As illustrated in Figure 2, the annotation ”Abnormal” signifies that the image deviates from normalcy, while the descriptor ”Severe Diabetic Macular Edema” denotes the specific disease depicted in the image. The statement ”lots of hard exudates near the macula center observed, Moderate Non-Proliferative Diabetic Retinopathy, Retinal hemorrhages, or hard exudates observed” offers a detailed explanation of this disease. The inclusion of sign-level explanations is pivotal for accurate disease diagnosis. To innovate within this realm, we introduce corresponding sign-level labels to each image, namely Vascular, Macular, FBC (Fundus Boundary Color), OCD (Optic Cup Disc), FHE (Fundus Hemorrhages Exudation), and Other. Specifically, vascular pertains to ocular artery and vein-related conditions. Macular addresses diseases within or associated with the macular area. FBC involves signs affecting the boundary areas and the overall fundus background, such as the Leopard Fundus. OCD encompasses diseases related to the cup and disc. FHE focuses on diseases related to retinal hemorrhage and exudation. We generate the corresponding sign labels for each image using the original label information of fundus images based on above standards. The distinction sets MMFundus apart from other multimodal datasets.
III-C The architecture of VisionUnite